If I have 17.0 liters of gas at a temperature of 67.0 °C and a pressure of 88.89 atm, what will be the pressure of the gas if I raise the temperature to 94.0 °C and decrease the volume to 12.0 L?(round to correct sig figs)


1
Expert's answer
2021-08-23T02:56:23-0400

This is an example of the combined gas law. The equation to use is:

(P1V1)/T1=(P2V2)/T2,

where P is the pressure, V is the volume, and T is the temperature in Kelvins. For this question, the temperature in degrees Celsius will be converted to Kelvins by adding 273.15.

P1= 88.89 atm

V1=17 L

T1= 67∘C + 273.15 = 340 K

V2=12 L

T2=94∘C+273.15=367 K

Lets rearrange the equation to isolate P2, substitute the known values into the equation, and solve.

P2=(P1xV1xT2)/(V2xT1)

P2=(88.89 atm × 17 L × 367 K) / (12 L × 340 K) = 140 atm rounded to two significant figures